# Adaptive Reduced Tempering For Bayesian inverse problems and rare event simulation

### v1

@@ 4.4.3 Bayesian inversion (Problem #BI)
Figure 7: Problem #BI (Bayesian inversion) & model #RB (PDE-based score). Kolmogorov-Smirnov distance as a function of the expected cost for ART’s IS and RSMC estimates for a varying number of particles $N$. Comparison to the performance of adaptive tempering with $S^{\ast}$.
@@
In contrast to the case of problem #RB with model #S, we find that ART’s IS estimate performs slightly worse than ART’s RSMC estimate. Indeed, its slope is less than the slope minus one half observed for the RSMC estimate. In this case, we conjecture that for sufficiently large $N$ and a sufficiently accurate approximation $S$, the bias of the ART’s IS and RSMC normalized estimates are negligible in comparison to their variances, and furthermore, the variance of ART’s IS estimate will be greater than that of ART’s RSMC estimate. Indeed, let’s assume that $S^{(k)}\simeq S^{\ast}$ for iterations $k>K_{j_{0}}$ (which is a plausible assumption based on the illustrations in figures 2 and 3). The RSMC estimate should then have a variance of at best $1/{NH}$ [15, Prop. 9.5.6] and a negligible bias [17]. On the other hand, based on the Appendix A, the IS estimate should have a variance of the order of $1/{H}$ and a negligible bias. Consequently, considering $N$ tending to infinity and $H$ fixed, the mean square error of the RSMC should tend towards zero with a speed at best inversely proportional to the cost (4.5) (of the order of $KN$), while the mean square error of the IS estimate will tend at best towards a constant (of the order of $1/H$).
In summary, our numerical experiments have demonstrated that a computational saving of over a decade is achieved with the ART algorithm, compared with state-of-the-art standard approaches. In addition, the product of variance and cost appears to be constant for these estimates in some regimes, suggesting asymptotic normality. The results also show that, for a particular problem, the RMSC estimate can be biased even in the case where the reduced score tends towards the true score. Conversely, on the Bayesian inverse problem, the RSMC estimate seems to perform slightly better than the IS estimate for a fixed budget of snapshots.
@@ 5 Conclusions and Perspectives
This work proposes an adaptive algorithm called ART that reduces the cost of adaptive SMC simulation when the score function $S^{\ast}$ is very expensive to compute. Riemann
The general idea of ART is to perform iterative importance sampling of a sequence of target distributions of the form $\propto e^{\beta^{(k)}{S^{\ast}}}d\pi$ parametrized at iteration $k$ by the inverse temperature $\beta^{(k)}\in\mathbb{R}_{+}$, using a sequences of proposals of the form $\mu_{\beta^{(k)}}^{(k)}\propto e^{\beta^{(k)}{S^{(k)}}}d\pi$, much cheaper to simulate with SMC. As the algorithm is iterated, the proposals are adapted by adjusting the inverse temperature $\beta^{(k)}$ and refining the approximation of the reduced score $S^{(k)}$ using the history of score evaluations previously calculated for importance sampling.
@@
In addition to demonstrating their consistency in an idealized setting, we evaluate in our numerical experiments the empirical convergence of the two proposed estimators, an importance sampling estimator and a reduced SMC estimator, in the context of rare-event simulation and for a Bayesian inverse problem based on a PDE. We exhibit empirically that we significantly reduce simulation cost of order of a decade while achieving a comparable accuracy.

### v2

@@ 4.4.3 Bayesian inversion (Problem #BI)
Fig. 8: Problem #BI (Bayesian inversion) & model #RB (PDE-based score). Kolmogorov-Smirnov distance as a function of the expected cost for ART’s IS and RSMC estimates for a varying number of particles NN. Comparison to the performance of adaptive tempering with $S^{\ast}$.
@@
In contrast to the case of problem #RB with model #S, we find that ART’s IS estimate performs slightly worse than ART’s RSMC estimate. Indeed, its slope is less than the slope minus one half observed for the RSMC estimate. In this case, we conjecture that for sufficiently large $N$ and a sufficiently accurate approximation $S$, the bias of the ART’s IS and RSMC normalized estimates are negligible in comparison to their variances, and furthermore, the variance of ART’s IS estimate will be greater than that of ART’s RSMC estimate. Indeed, let’s assume that $S^{(k)}\simeq S^{\ast}$ for iterations $k>K_{j_{0}}$ (which is a plausible assumption based on the illustrations in Figures 3 and 4). The RSMC estimate should then have a variance of at best 1/NH1/{NH} [18, Prop. 9.5.6] and a negligible bias [20]. On the other hand, based on the Appendix A, the IS estimate should have a variance of the order of 1/H1/{H} and a negligible bias. Consequently, considering $N$ tending to infinity and $H$ fixed, the mean square error of the RSMC should tend towards zero with a speed at best inversely proportional to the cost (4.6) (of the order of $KN$), while the mean square error of the IS estimate will tend at best towards a constant (of the order of $1/H$).
4.4.4 Model #RB in higher dimension
The core of the ART algorithm relies on sequential Monte Carlo methods for the tempering and bridging routines. Such methods are known to be stable in high dimensions [4]. However, in the specific case of Model #RB, the computational cost of ART increases substantially with the parameter space dimension $d$. More precisely, updating RB error approximations or evaluating this RB error induces a quadratic computational cost in $d$, inherent to the RB model and not the ART algorithm itself. It should be noted that the cost increase for update may be acceptable in the RB formalism as the reduced basis assembly is intended to be performed off-line. Consequently, as shown in Appendix B, increasing $d$ from $4$ to $25$ impacts ART performance with a speed loss of about two decades for a given expected square error, making the method non-competitive. We should nevertheless mention that there are possible remedies to avoid this slowdown with respect to $d$. Indeed, to avoid excessive cost of RB updates not affordable on-line, recent works challenge the RB off-line/on-line paradigm. In particular, the adaptive reduced modeling scheme of [37] proposes inexpensive on-line updates to the RB basis while maintaining its dimension. Other popular works on active subspace offer ways to circumvent the problem by learning a low-dimensional representation of the parameter space [16].
4.4.5 Summary
In summary, our numerical experiments have demonstrated that a computational cost reduction of over a decade is achieved with the ART algorithm, compared with state-of-the-art standard approaches. In addition, the product of variance and cost appears to be constant for these estimates in some regimes, suggesting asymptotic normality.
The results also show that, for a particular problem (Problem #S), the RMSC estimate can be biased even in the case where the reduced score tends towards the true score. Conversely, on the Bayesian inverse problem, the RSMC estimate seems to perform slightly better than the IS estimate for a fixed budget of snapshots.
@@ 5 Conclusions and Perspectives
This work proposes an adaptive algorithm called ART that reduces the cost of adaptive SMC simulation when the score function $S^{\ast}$ is very expensive to compute.
The general idea of ART is to perform iterative importance sampling of a sequence of target distributions of the form $\propto e^{\beta^{(k)}{S^{\ast}}}d\pi$ parametrized at iteration $k$ by the inverse temperature $\beta^{(k)}\in\mathbb{R}_{+}$, using a sequence of proposals of the form $\mu_{\beta^{(k)}}^{(k)}\propto e^{\beta^{(k)}{S^{(k)}}}d\pi$, much cheaper to simulate with SMC. As the algorithm is iterated, the proposals are adapted by adjusting the inverse temperature $\beta^{(k)}$ and refining the approximation of the reduced score $S^{(k)}$ using the history of score evaluations previously calculated for importance sampling.
@@
In addition to demonstrating their consistency in an idealized setting, we evaluate in our numerical experiments the empirical convergence of the two proposed estimators, an importance sampling estimator and a reduced SMC estimator, in the context of rare event simulation and for a Bayesian inverse problem based on a PDE. We exhibit empirically that we significantly reduce simulation cost of order of a decade while achieving a comparable accuracy.
@@ Appendix A Consistency
Appendix B ART & RB in high dimensions
We show in this appendix that increasing the parameter space dimension $d$ of the RB model (while keeping the other parameters in an analogous setting, see the table in Section 4.4.1), yields non-competitive results. The main reason for this performance drop is a quadratic complexity with respect to $d$: an RB update (with error) scales in $\mathcal{O}(hd^{2}K^{2})$), and an error evaluation scales as $\mathcal{O}(d^{2}K^{2})$. As a consequence, the following table
[…]
$S(x)$ v.s. $S^{\ast}(x)$ $S(x)$ & E(x) v.s. $S^{\ast}(x)$ ART v.s. SMC ART v.s. SMC
dd=4 $\times 63$ $\times 29$ $\times 17$ $\times 27$
dd=25 $\times 12$ $\times 3$ $\times 1$ $\times 2$
shows that increasing $d$ slows the ART algorithm and thus deteriorates its performance: while for $d=4$ the speed of the ART algorithm is for a given expected square error (4.4) nearly 20 times faster than adaptive SMC in terms of CPU time and more than 20 times less expensive in terms of the cost (4.6), both algorithms share comparable performance for $d=25$. A closer analysis of the previous table shows that increasing $d$ slows the relative cost (as compared to a true model evaluation) of solving an RB system (scaling in $\mathcal{O}(K^{3}+dK^{2})$) by a factor of 6, and adding the cost of error estimation (scaling in $\mathcal{O}(d^{2}K^{2})$) lowers this speed by up to a factor 10. More importantly, for $d=25$, the cost of RB updates becomes critically high even if rare, and exceeds the cost of the many evaluations of the RB solution and its error or the cost of computing snapshots, as detailed in the next table:
RB update tempering bridging $S^{\ast}$ eval.
%CPU time, dd=4 3.3% 63.7% (2.5% optim.) 32.0% (14.9% optim.) 1.0%
%CPU time, dd=25 68.4% 15.4% (<<0.1% optim.) 16.1% (0.4% optim.) 0.1%
We stress that the performance drop is due the computational complexity of the RB update and evaluation, especially of the error and does not come from the ART algorithm methodology itself.